OpenStudy (anonymous):

\[\int\limits_{1}^{4} (x^3+2x^2+16x+16)/(x^4+16x^2)*dx\]

OpenStudy (anonymous):

Solve, is there a semi-easy way to do this - without the PF that is.

OpenStudy (anonymous):

maybe a trick to write the numerator as the derivative of the denominator? would still have to split it up i think

OpenStudy (experimentx):

fist try to factorise .. that might give clue

OpenStudy (anonymous):

just thinking out loud but maybe \[\frac{x^3+16x}{x^4+16x^2}+\frac{2x^2+16}{x^4+16x^2}\] would that make it easier?

OpenStudy (anonymous):

\[x(x ^{2}+16) +2x ^{2}+16/x ^{2}(x ^{2}+16)\]

OpenStudy (anonymous):

no not really, but i bet we could break it up some way

myininaya (myininaya):

Does PF= partial fractions?

OpenStudy (anonymous):

I broke up it up into 4 different terms, and then began finding the PF decomposition of everyone of the terms.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

It's under that chapter in my book aswell. (Partial fractions that is)

OpenStudy (experimentx):

let's ask wolf ... we might construct solution.

OpenStudy (anonymous):

must be an easy trick since it is \[\frac{1}{x}+\frac{1}{x^2}+\frac{1}{x^2+16}\]

OpenStudy (experimentx):

if we walk in reverse direction

OpenStudy (anonymous):

easiest possible numerators, so probably some easy algebra will get it

OpenStudy (anonymous):

I got stuck on trying to decompose the last term :'(

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=integrate+%28x%5E3%2B2x%5E2%2B16x%2B16%29%2F%28+x%5E4%2B16x%5E2%29 looks like satellite73 was right after all.

OpenStudy (anonymous):

or just grind it out \[\frac{x^3+2x^2+16x+16}{x^4+16x^2}=\frac{a}{x}+\frac{b}{x^2}+\frac{cx+d}{x^2+16}\] \[x^3+2x^2+16x+16=ax(x^2+16)+b(x^2+16)+(cx+d)x^2\] put x = 0 get \(b=1\) right off the bat

OpenStudy (anonymous):

if we were snappy with complex numbers we could put x = 4i, but that is probaby too much

OpenStudy (anonymous):

but we have b = 1 so \(bx^2+dx^2=x^2+dx^2=2x^2\) now we know d = 1 as well

OpenStudy (anonymous):

and \(16x=16ax\) tells us \(a=1) also

OpenStudy (anonymous):

\(a=1\)

OpenStudy (anonymous):

Thanks alot! This is probably the most messy integral I've tried solving.

OpenStudy (anonymous):

And to go one direction - only to realize your way does not work in the last step - kills :(

OpenStudy (anonymous):

yeah these are a pain